OpenStudy (anonymous):

Find sin(x) if cos(x)=a/a+2.

OpenStudy (anonymous):

a) \[ \frac{ 2\sqrt{a+1} }{ a+2 }\] b) \[\frac{ 4(a+1) }{a+2 }\] c) \[\frac{ 2\sqrt{a+1} }{ a }\] d) \[\frac{ 4(a+1) }{ a }\] e) \[\frac{ \sqrt{a+1} }{2a }\]

OpenStudy (wach):

Try posting this in 'Mathematics'

thomaster (thomaster):

Post this one in the math section :P

OpenStudy (anonymous):

oh, opps

OpenStudy (anonymous):

No wonder... :P thx

OpenStudy (abb0t):

\[\cos(x)=\frac{ adjacent }{ hypotenuse }\]